OpenStudy (khally92):

Lagrange Multiplier Million Dollar question. Anybody please Use the method of Lagrange multipliers to find the max and min values of g(s,t)=t^2e^s given that s^2+t^2=3 . How are we assured that these extreme values exist?

OpenStudy (khally92):

I have one critical point

ganeshie8 (ganeshie8):

Are you looking for some real analysis kindof existence proof.. or just some intuition ?

OpenStudy (khally92):

real Analysis please

ganeshie8 (ganeshie8):

Then what is lagrange multiplier doing here ?

OpenStudy (khally92):

\[(1, \sqrt{2}) (1 -\sqrt{2})\]

OpenStudy (khally92):

Pardon?

ganeshie8 (ganeshie8):

Lagrange multiplier method is just one way to find the extrema. It has nothing to do with the existence part right ?

OpenStudy (khally92):

yes I know, and from the Extreme Value Theorem tells us that minimums and maximums will always exist.

OpenStudy (khally92):

I think my Instructor wants me to graph, and show that the max and min exist.

OpenStudy (khally92):

But I am only getting max and no mean from the two critical point \[(1 \sqrt{2}) (1 \sqrt{-2})\] and f gives me 2e for both points

OpenStudy (khally92):

Is this max or min point? How can I tell.

OpenStudy (khally92):

tried to graph but looks tedious.

OpenStudy (khally92):

and s also =-3 which has not value for t.

OpenStudy (khally92):

hello.

ganeshie8 (ganeshie8):

Lagrange multiplier doesn't give you information about the type of extrema. You will need to use hessian matrix for that

OpenStudy (khally92):

Hessian? did not cover that

ganeshie8 (ganeshie8):

Then you will need to find the type of extrema by some other means

OpenStudy (khally92):

but To determine if we have maximums or minimums we just need to plug these into the function. critical points

ganeshie8 (ganeshie8):

Yes that will do

OpenStudy (khally92):

yes that that is what I did.

OpenStudy (khally92):

Can you determine the max and min please?

OpenStudy (khally92):

I got 2e for boths critical points

OpenStudy (khally92):

Are the two points bot max? what about min. could it be that these function doest not have a min.

OpenStudy (khally92):

please look closely at x=-3 the value for t -root

ganeshie8 (ganeshie8):

\((\pm \sqrt{3}, 0 )\) give you minimum \((1,\pm \sqrt{2})\) give you maximum and yeah other two critical points are not real

ganeshie8 (ganeshie8):

I figured that simply by evaluating g(s,t) at those values, as you said earlier

OpenStudy (khally92):

And the last part of the question how are we sure that these values exist?

OpenStudy (khally92):

How did you get the second critical point pmsruareroont of 3.

ganeshie8 (ganeshie8):

\(\sqrt{3} = 1.73205\ldots\)

OpenStudy (khally92):

geometrical look at this problem is helpful. z= f (x, y)= x + 2y is a plane passing through the origin. The constraint x^2 +y ^2 + 5 is a right circular cylinder. The maximum value of f (x, y)= x+ 2y exists and occurs at the highest point on the curve of intersection of the cylinder and the plane. Example from my note. I do not need proof of lagrange to know if the values exist or not. i need like a geometrical look. @ganeshie8 and thank you so much. really apprecaite.

OpenStudy (khally92):

geometric look?

OpenStudy (khally92):

http://www.wolframalpha.com/input/?i=graph+t%5E2e%5Es form here I can see the two values of the max and min How can I plot this on paper. @ganeshie8 thank you.

OpenStudy (khally92):

@mathmale can you please look at this lagrange multiplier example. I get everything except the last part of the question. "How are we sure these extreme values exist" I know by the extreme value theorem a max and min WILL ALWAYS EXIST. Thank you

OpenStudy (khally92):

What will be the argument here? I could not sketch of the graph of function but from wolfram I can see that the function has two max value and one minimum value. http://www.wolframalpha.com/input/?i=graph+t%5E2e%5Es @mathmale Thanks

OpenStudy (khally92):

@mathmale help please.

OpenStudy (mathmale):

I'd love to help, but it's been several years since I last encountered Lagrange Multipliers. Frankly, I don't remember enough about them to help you. Thanks for calling on me, however. Good luck!

OpenStudy (khally92):

Its Ok Thanks